How many pounds of hamburger that costs $1.10 per pound must be mixed with 50 pounds of hamburger that costs $1.80 per pound to make a mixture that costs $1.30 per pound?
I came up with the answer of 175 pounds but I don't know how to write an equation to come up with this number.

Answer by scott8148(6628)

let x="lb @ $1.10" __ x(1.10)+50(1.80)=(x+50)(1.30) __ multiplying by 10 to clear decimals

11x+900=13x+650 __ 250=2x __ 125=x

175 is the total weight, not the weight of $1.10 hamburger